OpenStudy (anonymous):

10 + (10 – 12 ÷ 4 x 3) x 7 =

OpenStudy (anonymous):

First solve for brackets.. In brackets you will first divide, then multiply and subtraction.. Then multiply the result you got with 7 and then add 10 to it..

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Thanks for not just giving me the answer I'd like to solve it myself just wasn't sure how.

OpenStudy (anonymous):

Try it and if any doubt then tell me..

OpenStudy (anonymous):

I got 21.5, is that correct?

OpenStudy (anonymous):

I got 21.5, is that correct?

OpenStudy (anonymous):

Let me try..

OpenStudy (anonymous):

No check it once again..

OpenStudy (anonymous):

In brackets you can see \(-12 \div 4\) Solve this first..

OpenStudy (anonymous):

17?

OpenStudy (anonymous):

This is divide.. You can think of it as: \[\frac{-12}{4} = ??\]

OpenStudy (anonymous):

-12 / 4 =3

OpenStudy (anonymous):

Where - sign gone ??

OpenStudy (anonymous):

it = -3

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

So you are left with : \[10 + (10 \color{blue}{- 3} \times 3) \times 7\] Blue part is your result you got by solving -4 /12.. Okay ??

OpenStudy (anonymous):

Oh sorry -12/4 i mean to say..

OpenStudy (anonymous):

10 - 3 x 3 = 1. 1 x 7 +10 = 17. Right?

OpenStudy (anonymous):

Wow.. Right..

OpenStudy (anonymous):

Thank you I think I get it now.

OpenStudy (anonymous):

Welcome dear.. Try more you will definitely get those..